OpenStudy (anonymous):

factor 1(2x^2 - 5x - 2)

OpenStudy (anonymous):

what is the 1 doing out front?

OpenStudy (anonymous):

you have to find the GCF first then factor.. that was the directions.. its just 1..

OpenStudy (anonymous):

any help?????

OpenStudy (anonymous):

woops typo 12x^2 - 5x - 2

OpenStudy (alfie):

(3x-2)(4x+1)

OpenStudy (anonymous):

there is no such thing as the "greatest common factor" one thing. "common" means "in common" so you need at least two

OpenStudy (anonymous):

if you are supposed to write the polynomial as a produce, the directions should be "factor"

OpenStudy (anonymous):

.... the directions say check for GCF then factor each trinomial completely..

OpenStudy (anonymous):

can you just show me how to do it... ?

OpenStudy (anonymous):

i dont need an argument i came here for help

OpenStudy (anonymous):

what it means is check for the greatest common factor of each term. in this case it is 1, so you are just supposed to factor

OpenStudy (anonymous):

I said it was one... and i do not know how to factor it that is why i am asking for your help

OpenStudy (anonymous):

they wrote that in case you saw \[4x^2+8x+4\] and then you would start with \[4(x^2+2x+1)\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I see the answer, but I dont know how you got it.

OpenStudy (anonymous):

you can "factor by grouping" i suppose. i will show the steps for this one. but most people use "trial and error"

OpenStudy (anonymous):

the leading coefficient is 12, the constant is -2 multiply to get \[12\times -2=-24\] and now you have to think of two numbers whose product is -24 that add to -5

OpenStudy (anonymous):

-8 and 3 will work because \[-8\times 3=-24,-8+3=-5\]

OpenStudy (anonymous):

then split the middle term as \[12x^2+3x-8x-2\] and then factor each pair. you get \[3x(4x+1)-2(4x+1)\] and so you finish with \[(3x-2)(4x+1)\]

OpenStudy (anonymous):

Oh okay, that makes more sense.. thanks